# Detecting and Mitigating Insertion Hallucination in Video-to-Audio Generation

### v3

@@ 5.2 Validation of IH Metrics
We validated the reliability of our IH detection pipeline on a dedicated human-annotated dataset, where clips were manually labeled by consensus for the presence of speech and music hallucinations (see Appendix 7.1). However, the task of defining precise temporal boundaries for acoustic events is inherently subjective and prone to ambiguity. This makes evaluating a detector solely on its ability to perfectly replicate these boundaries (i.e., maximizing recall) a potentially misleading measure of performance. We contend that a more robust criterion for reliability is ensuring that any detected segment corresponds to a genuine, human-verified event, which places a stronger emphasis on precision.
To formally reflect this principle in our evaluation, we adopt the generalized $F_{\beta}$-score, defined as:
$F_{\beta}=(1+\beta^{2})\cdot\frac{\text{Precision}\cdot\text{Recall}}{\beta^{2}\cdot\text{Precision}+\text{Recall}}$
We specifically set $\beta=0.5$, which gives twice the weight to precision over recall, thereby aligning our quantitative evaluation with the goal of ensuring high detector reliability.
Table 1 reports the performance of our detection methods. Notably, the individual PANNs model achieves the highest F0.5 score (82.97). However, the Majority Vote (MV) ensemble attains a nearly identical score (82.62) while providing the added robustness of a multi-detector consensus. We therefore select MV for all subsequent experiments, concluding that the critical benefit of ensemble robustness outweighs the negligible difference in performance scores.
Table 1: Performance of individual detectors and fusion strategies on the human-annotated validation set, using the F0.5 score to emphasize precision. Best results are in bold, second-best are underlined.
Method Precision Recall F0.5 IoU
Individual Detectors
inaSpeechSegmenter 79.53 68.18 76.97 58.00
YAMNet 79.15 74.54 78.18 62.30
PANNs 85.40 74.51 82.97 66.09
Fusion Strategies
AND 90.93 55.45 80.61 52.54
OR 73.68 87.32 76.06 66.56
MV 84.94 74.49 82.62 65.80
@@ 5.3 Insertion Hallucination Assessment
We next apply our validated IH metrics to state-of-the-art V2A models and to evaluate the effectiveness of our proposed PFC method.
@@
Posterior Feature Correction substantially reduces hallucination. Across all benchmarks, PFC consistently lowers hallucination rates. The effect is strongest on the more diverse Kling-Audio-Eval and AVE datasets, where IH@vid and IH@dur drop by 40 to 65%. On VGGSound, the main training-domain dataset, PFC still reduces hallucination frequency by 45 to 52%, but the reduction in duration is smaller (10 to 25%). This indicates that models overfit to training biases, making in-domain hallucinations harder to suppress, and highlights PFC’s strength in improving generalization to out-of-domain data.
Conventional metrics remain robust, showing no systematic degradation. Crucially, this targeted reduction in hallucination does not come at the cost of overall generation quality. As shown in the $\Delta$ rows, conventional metrics such as FD, KL, and DeSync exhibit only minor fluctuations, with most changes falling below 3%. We even observe several instances of notable improvement, such as a 10.3% enhancement in FD and an 8.4% gain in KL for ThinkSound, suggesting that removing misleading visual features can sometimes help the model produce higher-quality audio. This stability confirms that PFC is a non-destructive method that precisely targets unwanted content without degrading audio quality, diversity, or temporal alignment.
Table 2: Results on Kling-Audio-Eval, VGGSound, and AVE, showing that PFC consistently reduces hallucinations without degrading quality or synchronization.
IH@vid $\downarrow$
IH@dur $\downarrow$
FD $\downarrow$
KL $\downarrow$
ISC $\uparrow$
IB $\uparrow$
DeSync $\downarrow$
Kling-Audio-Eval
GT 5.23 1.42
mmaudio 12.91 4.55 10.476 2.502 8.342 0.3425 0.6102
+ PFC 6.14 2.45 10.964 2.457 8.231 0.3420 0.6070
$\Delta$ 52.4% 46.2% 4.7% 1.8% 1.3% 0.1% 0.5%
thinksound 24.34 14.69 12.478 2.760 5.49 0.2067 0.7366
+ PFC 9.24 5.18 12.352 2.528 5.431 0.2174 0.7064
$\Delta$ 62.0% 64.7% 1.0% 8.4% 1.1% 5.2% 4.1%
VGGSound
GT 11.00 2.60
mmaudio 16.33 6.09 6.870 1.813 7.008 0.3391 0.6010
+ PFC 8.91 5.47 6.489 1.780 7.127 0.3384 0.5941
$\Delta$ 45.4% 10.2% 5.5% 1.8% 1.7% 0.2% 1.1%
thinksound 13.04 5.17 6.666 2.015 5.726 0.2263 0.7167
+ PFC 6.28 3.86 6.568 1.953 5.797 0.2242 0.7240
$\Delta$ 51.9% 25.4% 1.5% 3.1% 1.2% 0.9% 1.0%
AVE
GT 15.29 1.57
mmaudio 13.02 3.07 3.209 1.473 6.485 0.3776 0.5460
+ PFC 6.05 1.82 3.244 1.462 6.517 0.3771 0.5619
$\Delta$ 53.5% 40.7% 1.1% 0.7% 0.5% 0.1% 2.9%
thinksound 19.07 7.40 8.231 1.921 5.430 0.2500 0.7195
+ PFC 10.23 3.02 7.385 1.946 5.358 0.2505 0.7181
$\Delta$ 46.4% 59.2% 10.3% 1.3% 1.3% 0.4% 0.2%
5.4 Ablation Study of Replacement Strategies
@@
In contrast, our proposed PFC method consistently achieves the lowest IH@dur across all datasets, with average performance substantially surpassing all alternatives. These findings demonstrate that PFC’s effectiveness comes from its ability to precisely identify and correct the video segments most likely to mislead the model, rather than relying on arbitrary or random replacement.
5.5 Comparison with Alternative Correction Methods
Beyond ablating feature replacement strategies, we further compare PFC with alternative correction methods that exploit different modalities. To ensure robustness, we evaluate on the five Kling-Audio-Eval sublabels where the baseline exhibits the highest hallucination rates.
Results in Figure 4 (Right) show that PFC again achieves the best average performance, attaining the lowest mean IH@dur among all methods. This highlights its effectiveness even on the most challenging cases.
However, PFC is not universally superior. For object-centric categories such as Glass and Blender, detailed captions outperform PFC in reducing hallucinations, suggesting that fine-grained textual descriptions provide stronger semantic constraints than feature correction alone.
This complementary behavior points to a promising avenue for future research: combining input-level textual guidance (e.g., detailed captions or prompts) with our posterior, feature-level correction (PFC) could potentially yield an even more powerful and reliable method for suppressing hallucinations across diverse categories.
Figure 4: (Left) Ablation on feature correction, showing PFC consistently achieves the lowest IH@dur across datasets. (Right) Comparison with alternative correction methods on top-5 sublabels; PFC attains the best average IH@dur while detailed captions can outperform on object-centric classes.
@@ 6 Conclusion
We studied a neglected failure mode in video-to-audio generation: Insertion Hallucination (IH), where models synthesize speech or music that does not exist in the video. We introduced an evaluation framework that explicitly measures authenticity via a validated multi-detector ensemble and two metrics, IH@vid and IH@dur. On three benchmarks, these metrics reveal that state-of-the-art systems frequently hallucinate, even when conventional semantic and temporal scores appear strong.
To mitigate IH, we proposed Posterior Feature Correction (PFC), a training-free, two-pass procedure that detects hallucinated segments and masks the corresponding video features with a learned empty token in a second pass. PFC consistently reduces both the prevalence and duration of hallucinations across datasets while preserving (and sometimes improving) standard distribution, semantic, and synchronization metrics.
Ethics Statement
All experiments in this paper are conducted on publicly available datasets (VGGSound and Kling-Audio-Eval). The human-annotated validation set was created solely for evaluation, with no personally identifiable or sensitive information involved. Annotators participated voluntarily and were fairly compensated. Our method aims to improve model reliability by reducing hallucination, but we acknowledge that generative audio may still pose risks of misuse, which require broader safeguards beyond the scope of this work.
Reproducibility Statement
We describe datasets, models, detectors, and evaluation metrics in detail in Section 5. Appendix 7.1 provides the annotation pipeline. All code, configuration files, and annotated evaluation splits will be released to enable full reproduction of our results.

### v5

@@ 5.2 Validation of IH Metrics
We validated the reliability of our IH detection pipeline on a dedicated human annotated dataset, where clips were manually labeled for the presence of speech and music hallucinations (see Appendix). However, human annotators are generally insensitive to fine-grained acoustic boundaries and often overlook short pauses on the order of tens or even hundreds of milliseconds. As a result, relying solely on recall oriented evaluation can be misleading. A more reliable criterion is ensuring that each detected segment corresponds to a genuine human verified event, which calls for placing stronger emphasis on precision.
To quantify this emphasis, we adopt the generalized $F_{\beta}$-score:
$F_{\beta}=(1+\beta^{2})\cdot\frac{\text{Precision}\cdot\text{Recall}}{\beta^{2}\cdot\text{Precision}+\text{Recall}}.$
Precision oriented evaluation corresponds to $\beta<1$, though choosing $\beta$ too small would overly penalize recall and suppress meaningful detections. As shown in Fig. 4(b), the Majority Vote (MV) fusion strategy consistently achieves the highest $F_{\beta}$ scores across the range $\beta\in[0.40,0.92]$, outperforming both AND and OR. This interval also aligns with our goal of favoring precision without overcommitting to it, as it emphasizes precision while still preserving adequate recall.
The strong and stable performance of MV across this precision leaning interval indicates that it offers the most reliable trade-off between false positives and false negatives. We therefore adopt MV as the fusion strategy for all subsequent experiments.
Table 1: Results on Kling-Audio-Eval, VGGSound, and AVE, showing that HALCON consistently reduces hallucinations without degrading quality or synchronization.
Common Use Metrics Hallucination Metrics
Method FD $\downarrow$ KL $\downarrow$ ISC $\uparrow$ IB $\uparrow$ DeSync $\downarrow$ IH@vid $\downarrow$ IH@dur $\downarrow$
Kling-Audio-Eval
GT 5.2 1.4
MMAudio [cheng2025mmaudio] 10.48 2.50 8.34 0.34 0.61 12.9 4.6
+ HALCON 10.96 2.46 8.23 0.34 0.61 6.1  $\downarrow$ 52.4% 2.5  $\downarrow$ 46.2%
ThinkSound [liu2025thinksound] 12.48 2.76 5.49 0.21 0.74 24.3 14.7
+ HALCON 12.35 2.53 5.43 0.22 0.71 9.2  $\downarrow$ 62.0% 5.2  $\downarrow$ 64.7%
VGGSound
GT 11.0 2.6
MMAudio [cheng2025mmaudio] 6.87 1.81 7.01 0.34 0.60 16.3 6.1
+ HALCON 6.49 1.78 7.13 0.34 0.59 8.9  $\downarrow$ 45.4% 5.5  $\downarrow$ 10.2%
ThinkSound [liu2025thinksound] 6.67 2.02 5.73 0.23 0.72 13.0 5.2
+ HALCON 6.57 1.95 5.80 0.22 0.72 6.3  $\downarrow$ 51.9% 3.9  $\downarrow$ 25.4%
AVE
GT 15.3 1.6
MMAudio [cheng2025mmaudio] 3.21 1.47 6.49 0.38 0.55 13.0 3.1
+ HALCON 3.24 1.46 6.52 0.38 0.56 6.1  $\downarrow$ 53.5% 1.8  $\downarrow$ 40.7%
ThinkSound [liu2025thinksound] 8.23 1.92 5.43 0.25 0.72 19.1 7.4
+ HALCON 7.39 1.95 5.36 0.25 0.72 10.2  $\downarrow$ 46.4% 3.0  $\downarrow$ 59.2%
@@ 5.3 Insertion Hallucination Assessment
We next apply our validated IH metrics to state-of-the-art V2A models and to evaluate the effectiveness of our proposed HALCON method.
@@
HALCON substantially reduces hallucination. Across all benchmarks, HALCON consistently lowers hallucination rates. The effect is strongest on the more diverse Kling-Audio-Eval and AVE datasets, where IH@vid and IH@dur drop by 40 to 65%. On VGGSound, the main training-domain dataset, HALCON still reduces hallucination frequency by 45 to 52%, but the reduction in duration is smaller (10 to 25%). This indicates that models overfit to training biases, making in-domain hallucinations harder to suppress, and highlights HALCON’s strength in improving generalization to out-of-domain data.
Conventional metrics remain robust, showing no systematic degradation. Crucially, this targeted reduction in hallucination does not come at the cost of overall generation quality. As shown in the $\Delta$ rows, conventional metrics such as FD, KL, and DeSync exhibit only minor fluctuations, with most changes falling below 3%. We even observe several instances of notable improvement, such as a 10.3% enhancement in FD and an 8.4% gain in KL for ThinkSound, suggesting that removing misleading visual features can sometimes help the model produce higher-quality audio. This stability confirms that HALCON is a non-destructive method that precisely targets unwanted content without degrading audio quality, diversity, or temporal alignment.
5.4 Efficient Inference for HALCON
The standard HALCON pipeline consists of three stages: an initial audio generation in Stage 1, a fast hallucination detection and feature correction in Stage 2, and a full regeneration in Stage 3. This three-stage generation design introduces substantial inference overhead, with Stage 1 being the primary computational bottleneck.
We therefore examine whether Stage 1 truly requires a high-fidelity generation. Our hypothesis is that a lower-fidelity audio—obtained with fewer sampling steps—may still reveal hallucination artifacts clearly enough for effective detection.
To test this, we vary the number of sampling steps $N$ used in Stage 1 on a subset of Kling-Audio-Eval, while keeping the Stage 3 generation fixed at 24 steps. After an early-terminated Stage 1, we immediately perform detection, correction, and a full 24-step regeneration.
Figure 5 reports the results: the blue curve shows the final IH@dur, and the red curve shows total inference time. Two observations emerge. First, small $N$ leads to low-fidelity audio that harms detection and yields high IH@dur, but performance rapidly improves and stabilizes as $N$ increases. Second, inference time grows almost linearly with $N$, with the full three-stage setting exceeding 1.8s.
The gray region highlights a favorable performance–efficiency trade-off. These findings show that HALCON’s detection stage is robust to generation fidelity: using only a fraction of the sampling steps in Stage 1 achieves nearly the same hallucination suppression while substantially reducing end-to-end inference time. This greatly improves the practicality of deploying HALCON in real systems.
Figure 5: IH@dur and inference time versus Stage 1 sampling steps, highlighting the trade-off.
5.5 Ablation of Feature Correction Strategies
@@
In contrast, HALCON consistently achieves the lowest IH@dur on all datasets, with average performance substantially surpassing all alternatives. These findings demonstrate that HALCON’s effectiveness arises not merely from feature replacement, but from its ability to precisely locate and correct the hallucination-triggering video segments. Arbitrary or misaligned replacements cannot achieve comparable gains.
Figure 6: Ablation of feature correction strategies. HALCON yields the lowest IH@dur.
5.6 Comparison with Alternatives
Beyond analyzing feature replacement strategies, we further compare HALCON with alternative hallucination reduction methods that operate at different modality levels. For robustness, all evaluations are conducted on the five Kling-Audio-Eval sublabels where the baseline model exhibits the highest hallucination rates. We report both the average IH@dur and the highest IH@dur observed within each setting, capturing overall hallucination tendency as well as worst case behavior.
As shown in Figure 7, HALCON achieves the lowest average IH@dur and substantially reduces the highest IH@dur among all evaluated methods. This demonstrates that HALCON effectively suppresses hallucinations not only on average but also in extreme failure cases, highlighting its robustness on the most challenging categories.
However, HALCON is not universally superior. In certain object-centric categories, detailed captions achieve stronger reductions in hallucination, suggesting that fine-grained textual descriptions provide semantic constraints that feature-level correction alone cannot supply.
These complementary behaviors indicate that HALCON and text-driven approaches address different aspects of hallucination. Combining input-level textual guidance (e.g., detailed captions) with our posterior, feature-level correction (HALCON) may yield an even more powerful and reliable method for suppressing hallucinations across diverse scenarios.
Figure 7: Comparison with alternative hallucination reduction methods. HALCON achieves the best IH@dur.
@@ 6 Conclusion
In this work, we investigate the problem of Insertion Hallucination (IH) in V2A generation cases where models produce audio events without any visual support and demonstrate that this phenomenon is both common and systematically overlooked by existing evaluation metrics. To quantify IH, we introduce an end-to-end evaluation framework that combines a multi-detector ensemble with two metrics, IH@vid and IH@dur, enabling reliable measurement of both the prevalence and severity of hallucinations across datasets.
To mitigate IH, we propose HALCON, a three-stage inference-time conditioning method that identifies hallucinated segments, suppresses their misleading visual cues, and regenerates audio using corrected features. Extensive experiments show that HALCON consistently reduces hallucination rates by a large margin while maintaining and in some cases improving distributional, semantic, and synchronization quality.
Our findings highlight IH as a fundamental reliability challenge for V2A systems and demonstrate that addressing hallucinations requires going beyond conventional alignment oriented metrics. We hope this work encourages future research on building more trustworthy, visually grounded audio generation models and developing evaluation protocols that better reflect real-world deployment needs.
7 Appendix
7.1 Hallucination Detection Pipeline
Overview
Our hallucination detection pipeline operates on fixed-length clips of $9.98$ s. Given an audio file, each detector produces a set of segments $\{(l,s,e)\}$ where $l\in\{\texttt{speech},\texttt{music}\}$ and $0\leq s<e\leq T$. For each clip we compute
$\text{IH@vid}=\mathbb{1}[d_{\mathrm{IH}}>0],\qquad\text{IH@dur}=d_{\mathrm{IH}}/T,$
where $d_{\mathrm{IH}}$ is the union duration of all speech/music segments and $T=9.98$ s is the clip length. Union duration is computed after merging overlapping speech/music intervals.
For robustness we use three off-the-shelf detectors: inaSpeechSegmenter, YAMNet and PANNs-CNN14. All detectors are run in CPU-only mode with restricted threading to avoid contention in multi-processing. For each detector we keep only speech/music segments and store them in CSV/JSON files for later fusion and analysis. Unless otherwise specified, we use the default weights released by the corresponding authors.
inaSpeechSegmenter (ISS)
We use the official inaSpeechSegmenter implementation with the following settings:
  • Engine: vad_engine=’smn’, detect_gender=False.
  • Classes: From the raw ISS output we keep only segments whose label is either speech or music.
  • Post-processing: We round start/end times to milliseconds and discard non-positive-length segments.
ISS directly outputs contiguous segments, so we do not change its internal thresholds; all later temporal smoothing is performed by our shared post-processing described below.
YAMNet
For YAMNet we use the TensorFlow Hub model google/yamnet/1 and compute frame-level scores at 16 kHz. Each frame corresponds to a 0.96 s receptive field with 50% overlap; we approximate the frame hop as $t_{\mathrm{hop}}=0.48$ s.
We map the original AudioSet labels to speech and music using curated label lists. Let $\mathcal{I}_{\mathrm{sp}}$ and $\mathcal{I}_{\mathrm{mu}}$ denote the indices of classes belonging to speech and music respectively. For each frame $i$ we compute
$p^{(i)}_{\mathrm{sp}}=\max_{c\in\mathcal{I}_{\mathrm{sp}}}p^{(i)}_{c},\qquad p^{(i)}_{\mathrm{mu}}=\max_{c\in\mathcal{I}_{\mathrm{mu}}}p^{(i)}_{c}$.
A frame is marked as speech if $p^{(i)}_{\mathrm{sp}}\geq 0.40$ and as music if $p^{(i)}_{\mathrm{mu}}\geq 0.30$. Each positive frame $i$ is converted to a tentative segment $s_{i},e_{i}]=[i\cdot t_{\mathrm{hop}},(i+1)\cdot t_{\mathrm{hop}}]$ with the corresponding label. These frame-level segments are then merged using our generic merging function (described below) with $\texttt{min\_dur}=0.20$ s and $\texttt{min\_gap}=0.15$ s.
PANNs (CNN14)
We use the SoundEventDetection interface from panns_inference with the CNN14 backbone pretrained on AudioSet. Audio is resampled to 32 kHz and fed as a single-channel waveform.
The model outputs a frame-wise tensor of shape $T\times C$. We normalise label names to lower case and map them into the same speech and music groups as in the main paper. Let $\mathcal{J}_{\mathrm{sp}}$ and $\mathcal{J}_{\mathrm{mu}}$ denote the corresponding index sets. For each time frame $i$ we compute
$q^{(i)}_{\mathrm{sp}}=\max_{c\in\mathcal{J}_{\mathrm{sp}}}q^{(i)}_{c},\qquad q^{(i)}_{\mathrm{mu}}=\max_{c\in\mathcal{J}_{\mathrm{mu}}}q^{(i)}_{c}$.
We mark frame $i$ as speech if $q^{(i)}_{\mathrm{sp}}\geq 0.40$ and as music if $q^{(i)}_{\mathrm{mu}}\geq 0.30$. Let $T$ be the number of frames and $L$ the clip duration; the corresponding frame hop is $t_{\mathrm{hop}}=L/T$. As with YAMNet, positive frames are converted into segments $s_{i},e_{i}]$ and merged by the common post-processing with $\texttt{min\_dur}=0.20$ s and $\texttt{min\_gap}=0.15$ s.
Temporal Smoothing and Merging
All detectors share the same temporal smoothing procedure. Given a list of frame-level segments $(s,e,l)$, we first sort them by start time and then iteratively merge consecutive segments with the same label if the gap between them is shorter than $\texttt{min\_gap}=0.15$ s. Segments shorter than $\texttt{min\_dur}=0.20$ s are discarded. This removes spurious short bursts and enforces temporal consistency across detectors.
Multi-detector Fusion
To combine the three detectors, we discretise the time axis into a uniform grid with step $\Delta t=0.02$ s. For each detector, we project its speech/music segments onto this grid (with a tolerance of 0.05 s) and obtain binary vote matrices of shape $K\times N$, where $K$ is the number of detectors and $N$ is the number of grid frames.
We consider three fusion rules:
  • OR: a frame is positive if at least one detector votes positive;
  • AND: a frame is positive only if all detectors agree;
  • MV (Majority Vote): a frame is positive if at least $\lceil K/2\rceil$ detectors vote positive.
Positive frames are then converted back to continuous segments and merged with the same $\texttt{min\_dur}=0.20$ s and $\texttt{min\_gap}=0.15$ s constraints. All main results in the paper use the MV fusion rule, which we show in Sec. 5.2 to offer the best precision–recall trade-off on our human-annotated validation set.
7.2 Human Annotation Pipeline and Use for Metric Validation
Sampling.
We construct a human-annotated validation set from the outputs of two representative V2A systems, MMAudio and ThinkSound, on the Kling-Audio-Eval benchmark. For each sublabel in Kling-Audio-Eval, we randomly sample $20$ generated audio clips per model, yielding over $900$ clips in total and more than 9,0009{,}000 seconds of audio. [^1] Counts refer to post-filtered items that pass loading and playback checks. This per-sublabel sampling strategy provides broad coverage of object- and scene-centric categories while keeping the annotation workload tractable.
Figure 8: Web interface for human annotation. Annotators can view the paired video and the spectrogram, select arbitrary time spans for playback, and mark multiple speech or music segments with $0.01$ s precision.
Annotation Interface.
Annotators use a web-based tool that supports precise temporal labeling (Figure 8). They can view the paired video, inspect the mel-spectrogram, scrub or play arbitrary time spans, and add multiple speech or music segments per clip. Timestamps are recorded at $0.01$ s resolution, and the interface enforces basic consistency constraints.
Labeling Guidelines.
Annotators mark a segment only if the generated audio contains audible speech or music without a plausible visual source. Ambiguous non-speech vocalisations are ignored unless clearly constituting speech or singing. Environmental noise and percussive impacts are not labeled. Segments shorter than $0.20$ s are discouraged, and gaps shorter than $0.15$ s between adjacent segments of the same type may be merged.
Boundary Conventions.
Boundaries align to perceptual onset/offset at $0.01$ s granularity. Adjacent labels separated by gaps <0.15<0.15 s are merged. During evaluation of detector accuracy, we apply a symmetric tolerance of $0.05$ s around annotated boundaries (Section 7.2).
Annotator Roles.
Two primary annotators independently label every clip, and a third adjudicator resolves disagreements. All annotators work with over-ear headphones in a quiet room, performing a brief warm-up on five clips with discussion of edge cases.
Quality Control.
Checks include schema validation (types, ordering), consistency validation (no impossible overlaps), and cross-annotator comparison (IoU and clip-level agreement). If disagreement persists, the adjudicator produces the final consensus annotation.
Aggregation and Export.
We export per-clip annotations with fields: clip_id, model, sublabel, segment_type, start, end. Timestamps are rounded to $0.01$ s. Post-processing merges gaps <0.15<0.15 s and removes fragments shorter than $0.20$ s unless manually retained. All annotations remain in the clip’s native time base.
Use in Validating IH Metrics.
For metric validation, we treat hallucination detection as a framewise binary classification problem on a fixed temporal grid. Both human annotations and detector outputs are converted into binary masks at a resolution of $0.01$ s over a 10 s clip. Adjudicated human segments are first clamped to $[0,10]$ and rasterized into a ground-truth mask $m^{\ast}$. For each detector (ina, panns, yamnet), we load its CSV file, select all rows associated with the current clip, clamp predicted timestamps, and rasterize them into a mask $m$ on the same grid. An optional tolerance $\tau$ can be applied via one-dimensional binary dilation, though the main results use $\tau=0$.
For each clip and each method (three base detectors and three fusion strategies: AND, OR, and majority vote), we compute per-frame true positives, false positives, and false negatives as $\mathrm{TP}=\sum(m^{\ast}\wedge m)$, $\mathrm{FP}=\sum(\neg m^{\ast}\wedge m)$, $\mathrm{FN}=\sum(m^{\ast}\wedge\neg m)$, and accumulate these counts over all annotated clips. We then compute precision, recall, and IoU, and evaluate the $F_{\beta}$ score over $\beta\in[0,1]$:
$F_{\beta}=(1+\beta^{2})\cdot\frac{\mathrm{Precision}\cdot\mathrm{Recall}}{\beta^{2}\,\mathrm{Precision}+\mathrm{Recall}}.$
Varying $\beta$ allows us to smoothly adjust the relative importance of precision and recall, enabling a more complete characterization of detector behavior. This evaluation pipeline ensures strict temporal alignment between human judgments and detector predictions and provides a consistent basis for comparing individual detectors and fusion strategies.
Statistics.
The final annotated set contains over 1,0001{,}000 clips (more than 10,00010{,}000 seconds). Speech and music prevalence varies substantially across sublabels. We report hallucination prevalence and average hallucinated duration in Section 5.2.
Intended Use.
This annotated set is used solely for evaluating the IH detection pipeline and is never used to train or tune HALCON. All quantitative results relying on human labels reference the adjudicated split.
7.3 Case Studies of Hallucinated Speech and Music
Case 1: Speech hallucination.
Figure 9 shows a representative example in which the video depicts a buffalo grazing in a field, a scene that contains no plausible speech source. Nevertheless, the audio produced by ThinkSound includes two hallucinated speech segments, whereas HALCON successfully suppresses both artifacts and remains consistent with the natural acoustic scene.
In the ThinkSound spectrogram, two speech-like regions emerge around 5.5–6.5 s and 8.2–9.0 s. The first region is faint: it contains several short, weak, and discontinuous high-frequency streaks that create only a vague vocal impression. The second region displays clear speech structure, including: (i) stable formant-like trajectories between 300–3,000 Hz, (ii) periodic temporal patterns resembling syllabic rhythm, and (iii) transient energy bursts indicative of consonant onsets. Listening confirms that the second segment corresponds to a clearly audible male voice saying “fifty nine,” which is unrelated to the visual scene and strongly disrupts perceptual coherence.
In contrast, HALCON exhibits no formant patterns, no periodic vocal structures, and no consonant-like transients throughout the clip. Its time–frequency texture remains stable and consistent with natural grazing sounds such as foliage rustling, low-frequency animal movement, and ambient outdoor noise. This matches the ground-truth (GT) spectrogram, which contains no speech activity.
Case 2: Long-duration music hallucination.
Figure 10 provides another example, where the video shows a train moving through a rural landscape. The true audio in such a scene is dominated by low-frequency rolling noise, wind, and occasional broadband transients, with no plausible source of music. However, ThinkSound generates a persistent music-like pattern that spans nearly the entire 0–10 s duration.
The hallucinated music is characterized by: (i) dense vertical striations that resemble synthetic percussive or rhythmic elements, (ii) quasi-periodic temporal repetition suggestive of a musical beat or harmonic cycle, and (iii) stable high-frequency energy bands inconsistent with environmental noise. These patterns collectively form a clearly perceptible “background music” layer that overrides the natural ambience expected from the scene.
HALCON suppresses these artifacts entirely. Its spectrogram exhibits smooth, broadband low-frequency energy consistent with train motion, sparse and irregular high-frequency components typical of wind and environmental texture, and no periodic or harmonic spectral structures. Compared with the GT spectrogram, HALCON closely preserves the authentic acoustic characteristics of the scene and avoids introducing any music-like artifacts.
Summary.
Across both speech and music examples, ThinkSound produces salient, perceptually dominant hallucinated content—either short but distinct speech or long-duration music. HALCON consistently removes such spurious vocal and musical artifacts while maintaining scene-appropriate acoustic structure, achieving outputs that are both perceptually and physically aligned with the visual input.
Figure 9: Case study showing hallucinated speech in ThinkSound. The video shows a buffalo grazing, a scene with no plausible speech source. ThinkSound introduces two spurious speech segments (indicated by green markers), while HALCON eliminates both and aligns closely with the ground-truth (GT) acoustic structure.
Figure 10: Case study showing hallucinated music in ThinkSound. The video records a train moving through a rural landscape, a scene with no plausible music source. ThinkSound introduces a long-duration music-like pattern that spans almost the entire clip, whereas HALCON suppresses these artifacts and produces an output that is consistent with the ground-truth (GT) ambient acoustic structure.